Covid-19 on copper metal is said to be virulent within 4 hours. Jung Hyuk wants to check if the metal piece he found in his father's drawer is indeed copper. He finds that the metal is 125g in air and 111g when suspended in water. Is this metal copper?
1
Expert's answer
2020-04-21T19:03:13-0400
"m=\rho gV, m_w=(\rho -\rho_w)gV"

"\frac{m_w}{m}=1-\frac{\rho_w}{\rho}\to \frac{111}{125}=1-\frac{1000}{\rho}"

"\rho=8930\frac{kg}{m^3}"

Yes, this is copper.